OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 24 ?

OpenStudy (anonymous):

a1 = 4 = 1*4 a6 = 24 = 6*4 a32 = 32*4 = 128

OpenStudy (anonymous):

a1=4 a6 = a1+5d=24 therefore d = 4 a32 = a1+31d=4+31(4) = 128